# BatteryLife: A Comprehensive Dataset and Benchmark for Battery Life Prediction

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; August 03–07, 2025; Toronto, ON, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data mining
@@ 1. INTRODUCTION
SNL (Preger et al., 2020) 50 1 3 3 23 Lab test Li-ion 2020
MICH_EXP (Mohtat et al., 2021) 18 1 1 3 6 Lab test Li-ion 2021
@@
BatteryML (Zhang et al., 2023a) 381 2 5 5 186 Lab test Li-ion 2024
@@
  • Limited dataset size. As shown in Table 1, the previous largest dataset is BatteryML (Zhang et al., 2023a) that offers 381 end-of-life batteries by integrating 7 public datasets (Attia et al., 2020; Severson et al., 2019; Ma et al., 2022; Devie et al., 2018; Preger et al., 2020; Juarez-Robles et al., 2020; Li et al., 2021; Juarez-Robles et al., 2021). However, this undertaking overlooks several datasets (Zhu et al., 2022; Weng et al., 2021; Mohtat et al., 2021) and lacks datasets (Cui et al., 2024; Wang et al., 2024c; Li et al., 2024) that are released recently. In this vein, BatteryML, whose size is smaller than half of the sum of existing datasets, fails to provide comprehensive insights into modern battery life data.
@@
To address these issues, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. From a dataset perspective, BatteryLife stands out as the largest and most diverse battery life dataset by integrating 16 datasets. Specifically, BatteryLife is 2.5 times the size of BatteryML (Zhang et al., 2023a) in terms of battery number, covering more than 90,000 samples. Remarkably, BatteryLife furnishes unparalleled diversity, delivering 4 times format, 11.8 times chemical system, 1.8 times operation temperature, and 2.2 times charge/discharge protocol compared to BatteryML (Zhang et al., 2023a). Notably, BatteryLife pioneers as the first work that releases zinc-ion (Zn-ion) and sodium-ion (Na-ion) battery life datasets, and additionally offers large-capacity Li-ion batteries that are tested by the battery manufacturer (i.e., CALB Group Co., Ltd.). The formidable diversity challenges the model’s capability to learn essential mappings from early data to life and allows inspiration for the data mining community and a broader battery community.
@@
  • The largest battery life dataset: BatteryLife offers more than 90,000 samples from 990 batteries, which is 2.5 times the size of the previous largest one. The unprecedented dataset size allows researchers to generate practical insights into battery life data by making advances on BatteryLife.
@@ 4.4. Ablation Studies (RQ2)
We conduct ablation studies of the proposed CyclePatch to study the effectiveness of its key components, and present the results in Table 4. Specifically, we study CPMLP and CPTransformer considering their leading performance across datasets. ”w/o Intra” refers to the removal of the intra-cycle encoder; ”w/o Inter” indicates the absence of the inter-cycle encoder; MLP and Transformer severs as naive baselines without patching. From Table 4, it is evident that all components in CyclePatch significantly contribute to its performance. Concretely, the incorporation of the intra-cycle encoder enhances the model’s ability to model interactions among variables within cycles, leading to performance gains across all datasets. Removing the inter-cycle encoder reduces performance, highlighting the importance of modeling inter-cycle interactions. Moreover, it is observed that CPMLP without either intra-encoder or inter-encoder performs worse than MLP on Na-ion and CALB datasets, whereas MLP naively models the intra-cycle and inter-cycle information. This indicates that modeling intra-cycle and inter-cycle information are both critical. These results collectively validate the effectiveness of the CyclePatch design.
@@ A.3. Overview of The Integrated Datasets
  • SNL (Preger et al., 2020): This dataset consists of 50 batteries in the format of 18650 cylindrical batteries. The positive electrode covers $\rm LiFePO_{4}$, $\rm LiNi_{0.81}Co_{0.14}Al_{0.05}O_{2}$, and $\rm LiNi_{0.84}Mn_{0.06}Co_{0.1}O_{2}$, while the negative electrode is graphite. The electrolyte formula is unknown. There are 3 nominal capacities consisting of 1.1 Ah, 3.2 Ah, and 3.0 Ah. There are 23 different protocols in this dataset with 3 different working temperatures (15°C, 25°C and 35°C).
@@
  • MICH_EXP (Mohtat et al., 2021): This dataset consists of 18 lithium-ion batteries in the format of pouch batteries. The positive electrode is $\rm LiNi_{1/3}Co_{1/3}Mn_{1/3}O_{2}$ (NCM111), and the negative electrode is $\rm graphite$. The electrolyte consists of 1M $\rm LiPF_{6}$ salt in solvents of ethylene carbonate (EC) and ethyl methyl carbonate (EMC) with a ratio of 3:7. The nominal capacity is 5.0 Ah. There are 6 different protocols with 3 different working temperatures in total (-5°C, 25°C, and 45°C).
@@
  • Zn-ion: This dataset consists of 100 zinc-ion batteries in the format of coin batteries. The positive electrode is $\rm MnO_{2}$ while the negative electrode is zinc metal. The electrolyte is unknown. The discharge capacity of the $10^{th}$ cycle was taken as the nominal capacity for each battery. All batteries underwent the same charge/discharge protocol at 3 different working temperatures.
@@ B.2. Implementation Details of Baselines
Regarding the CNN, we first reshape $\overline{\textbf{X}}_{1:100}$ into the shape of $\textbf{S}\in\mathbb{R}^{3,100,300}$, where 3 indicates the three variables (voltage, current and capacity), 100 is the number of maximum input cycles, 300 denotes the resampled length. Then, a CNN is utilized to model the S. In each layer, a Conv2d is first used, followed by ReLU activation. After the activation, average pooling is used to extract useful information. After a preset number of layers, a linear projection is used to produce the prediction.

### v4

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737372isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Data mining
@@ 1. INTRODUCTION
SNL (Preger et al., 2020) 52 1 3 3 23 Lab test Li-ion 2020
MICH_EXP (Mohtat et al., 2021) 12 1 1 3 6 Lab test Li-ion 2021
@@
BatteryML (Zhang et al., 2023a) 383 2 5 5 186 Lab test Li-ion 2024
@@
  • Limited dataset size. As shown in Table 1, the previous largest dataset is BatteryML (Zhang et al., 2023a) that offers 383 end-of-life batteries by integrating 7 public datasets (Attia et al., 2020; Severson et al., 2019; Ma et al., 2022; Devie et al., 2018; Preger et al., 2020; Juarez-Robles et al., 2020; Li et al., 2021; Juarez-Robles et al., 2021). However, this undertaking overlooks several datasets (Zhu et al., 2022; Weng et al., 2021; Mohtat et al., 2021) and lacks datasets (Cui et al., 2024; Wang et al., 2024c; Li et al., 2024) that are released recently. In this vein, BatteryML, whose size is smaller than half of the sum of existing datasets, fails to provide comprehensive insights into modern battery life data.
@@
To address these issues, we propose BatteryLife, a comprehensive dataset and benchmark for BLP. From a dataset perspective, BatteryLife stands out as the largest and most diverse battery life dataset by integrating 16 datasets. Specifically, BatteryLife is 2.5 times the size of BatteryML (Zhang et al., 2023a) in terms of battery number, covering 99,000 samples. Remarkably, BatteryLife furnishes unparalleled diversity, delivering 4 times format, 11.8 times chemical system, 1.8 times operation temperature, and 2.2 times charge/discharge protocol compared to BatteryML (Zhang et al., 2023a). Notably, BatteryLife pioneers as the first work that releases zinc-ion (Zn-ion) and sodium-ion (Na-ion) battery life datasets, and additionally offers large-capacity Li-ion batteries that are tested by the battery manufacturer (i.e., CALB Group Co., Ltd.). The formidable diversity challenges the model’s capability to learn essential mappings from early data to life and allows inspiration for the data mining community and a broader battery community.
@@
  • The largest battery life dataset: BatteryLife offers 99,000 samples from 990 batteries, which is 2.5 times the size of the previous largest one. The unprecedented dataset size allows researchers to generate practical insights into battery life data by making advances on BatteryLife.
@@ 4.4. Ablation Studies (RQ2)
We conduct ablation studies of the proposed CyclePatch to study the effectiveness of its key components, and present the results in Table 4. Specifically, we study CPMLP and CPTransformer considering their leading performance across datasets. ”w/o Intra” refers to the removal of the intra-cycle encoder; ”w/o Inter” indicates the absence of the inter-cycle encoder; MLP and Transformer serve as naive baselines without patching. From Table 4, it is evident that all components in CyclePatch significantly contribute to its performance. Concretely, the incorporation of the intra-cycle encoder enhances the model’s ability to model interactions among variables within cycles, leading to performance gains across all datasets. Removing the inter-cycle encoder reduces performance, highlighting the importance of modeling inter-cycle interactions. Moreover, it is observed that CPMLP without either intra-encoder or inter-encoder performs worse than MLP on Na-ion and CALB datasets, whereas MLP naively models the intra-cycle and inter-cycle information. This indicates that modeling intra-cycle and inter-cycle information are both critical. These results collectively validate the effectiveness of the CyclePatch design.
@@ A.3. Overview of The Integrated Datasets
  • SNL (Preger et al., 2020): This dataset consists of 52 batteries in the format of 18650 cylindrical batteries. The positive electrode covers $\rm LiFePO_{4}$, $\rm LiNi_{0.81}Co_{0.14}Al_{0.05}O_{2}$, and $\rm LiNi_{0.84}Mn_{0.06}Co_{0.1}O_{2}$, while the negative electrode is graphite. The electrolyte formula is unknown. There are 3 nominal capacities consisting of 1.1 Ah, 3.2 Ah, and 3.0 Ah. There are 23 different protocols in this dataset with 3 different working temperatures (15°C, 25°C and 35°C).
@@
  • MICH_EXP (Mohtat et al., 2021): This dataset consists of 12 lithium-ion batteries in the format of pouch batteries. The positive electrode is $\rm LiNi_{1/3}Co_{1/3}Mn_{1/3}O_{2}$ (NCM111), and the negative electrode is $\rm graphite$. The electrolyte consists of 1M $\rm LiPF_{6}$ salt in solvents of ethylene carbonate (EC) and ethyl methyl carbonate (EMC) with a ratio of 3:7. The nominal capacity is 5.0 Ah. There are 6 different protocols with 3 different working temperatures in total (-5°C, 25°C, and 45°C).
@@
  • Zn-ion: This dataset consists of 95 zinc-ion batteries in the format of coin batteries. The positive electrode is $\rm MnO_{2}$ while the negative electrode is zinc metal. The electrolyte is unknown. The discharge capacity of the $10^{th}$ cycle was taken as the nominal capacity for each battery. All batteries underwent the same charge/discharge protocol at 3 different working temperatures.
@@ B.2. Implementation Details of Baselines
Regarding the CNN, we first reshape $\overline{\textbf{X}}_{1:100}$ into the shape of $\textbf{S}\in\mathbb{R}^{3\times 100\times 300}$, where 3 indicates the three variables (voltage, current and capacity), 100 is the number of maximum input cycles, 300 denotes the resampled length. Then, a CNN is utilized to model the S. In each layer, a Conv2d is first used, followed by ReLU activation. After the activation, average pooling is used to extract useful information. After a preset number of layers, a linear projection is used to produce the prediction.